Grade 12th passGeneral Physics

a man can swim in still water with a speed of 3m/s. x and y axis are drawn along and normal to the bank of river flowing to right with the speed of 1 m/s .the man starts swimming from origin O at t=0 sec. find the equation of locus of all the possible points where man can reach at t=1 sec

mukul agrawal
10 Years agoGrade 12th pass

1 Answer

Tejeshwani Rathore
8 Years ago
Vx =3cosx.  Vnet x = 3cosx+1
Vy=3sinx 
X-1=3cosx
Y=3sinx
Squaring and adding.
(X-1)^2+y^2=9
(X-1)^2+y^2=3^2(cosx^2+sinx^2)